# Fan distributions via Tverberg partitions and Gale duality

### v2

@@ 5. Optimality of Fan Distributions and Robustness for Typical Point sets
In this section we show that there is a dense open class of affinely spanning point sets for which the distributing $r$-fans of Theorems 1.4 and 1.6 contain a robust fraction of the point set. In addition, we prove various cases of the near optimality of Theorem 1.4 with respect to dimension.
@@ 5.1. Typical Point Sets
Let $a_{1},\ldots,a_{n}$ be a sequence of $n$ affinely spanning points in $\mathbb{K}^{d}$ and let $a_{n+1}$ by the average of the $a_{i}$. Letting $A$ be the matrix of the Gale transform associated to $a_{1},\ldots,a_{n+1}$, one may choose the vector $(1,\ldots,1,-n)$ to be in the basis for $\ker A$. Thus the dual of $a_{1},\ldots,a_{n+1}$ is of the form $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=(-\sum_{i=1}^{n}x_{i},-n)$. We say that the sequence $x_{1},\ldots,x_{n}$ of affinely spanning points in $\mathbb{K}^{n-d-1}$ corresponds to $a_{1},\ldots,a_{n}$. It is not difficult to see that any sequence $x_{1},\ldots,x_{n}$ of affinely spanning points in $\mathbb{K}^{n-d-1}$ corresponds to some sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$. Moreover, it is an immediate consequence of Proposition 2.2 that a sequence of affinely spanning points $x_{1},\ldots,x_{n}$ in $\mathbb{K}^{n-d-1}$ which correspond to a sequence $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ are in general position (i.e., no $n-d-1$ of the $x_{i}$ lie on a hyperplane in $\mathbb{K}^{n-d-1}$) if and only if $a_{1},\ldots,a_{n}$ are in general position in $\mathbb{K}^{d}$ as well. For instance, if $x_{1},\ldots,x_{n-d-1}$, say, do not affinely span $\mathbb{K}^{n-d-1}$, then $g_{1},\ldots,g_{n-d-1},g_{n}$ lie on a linear hyperplane in $\mathbb{K}^{n-d}$ and so there is an affine dependence among $a_{n-d},\ldots,a_{n}$. The proof of the converse is nearly identical.
Our condition on affinely spanning point sets $X$ with $n$ elements in $\mathbb{K}^{n-d-1}$ is essentially the Gale dual of a generic condition on affinely spanning point sets $A$ of $n$ elements in $\mathbb{K}^{d}$. We recall that a sequence $a_{1},\ldots,a_{n}$ of $n$ points in $\mathbb{K}^{d}$ is in strong general position [23] if (1) the $a_{i}$ are in ordinary general position and (2) for any non-empty pairwise disjoint subsets $I_{1},\ldots,I_{r}$ of $[n]$, one has that the (real) codimension of the intersection of the affine hulls $\operatorname{\mathrm{Aff}}(A_{j})$ of the $A_{j}=\{a_{i}\mid i\in I_{j}\}$ equals the sum of their respective codimensions, that is,
@@
We remark that condition (1) follows from condition (2) when $\mathbb{K}=\mathbb{R}$ (see, [23]). The central property of strong general position for our purposes is that it immediately implies that any Tverberg $r$-tuple for a linear map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ determined by points $a_{1},\ldots,a_{n}$ in strong general position must satisfy $\sum_{j=1}^{r}|\operatorname{\mathrm{Vert}}(\sigma_{j})|\geq(r-1)(d+1)+1$.
We shall say that an affinely spanning point set $X=\{x_{1},\ldots,x_{n}\}$ in $\mathbb{K}^{n-d-1}$ is typical if $x_{1},\ldots,x_{n}$ corresponds to a sequence $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ in strong general position (the $x_{i}$ are in ordinary general position and are therefore distinct). While we expect that typical sets are generic, we shall content ourselves with showing that they form an open and dense in the space of all affinely spanning point sets of size $n$ in $\mathbb{K}^{n-d-1}$. To show this, we view any sequence of $n$ points in $\mathbb{K}^{d}$ as an element of $(\mathbb{K}^{d})^{n}$, and likewise for sequences of $n$ points in $\mathbb{K}^{n-d-1}$. Let $S$ be the set of all affinely-spanning sequences of length $n$ in $(\mathbb{K}^{d})^{n}$. By generiticity of strong general position, the subset $U$ of $S$ consisting of all such sequences which are in strong general position is a dense open subset of $S$. One has the analogous situation holds for typical point sets of size $n$ in $\mathbb{K}^{n-d-1}$.
@@
Let $S^{\prime}$ be the subset $(\mathbb{K}^{n-d-1})^{n}$ consisting of all affinely spanning sequences of length $n$ in $\mathbb{K}^{n-d-1}$ and let $U^{\prime}$ be the subset of $S$ consisting of all such sequences which correspond to a sequence of points in $\mathbb{K}^{d}$ in strong general position. Then $U^{\prime}$ is a dense open subset of $S^{\prime}$.
@@
We first prove density. Suppose that $x=(x_{1},\ldots,x_{n})$ is a sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$, so that $x$ corresponds to a sequence $a=(a_{1},\ldots,a_{n})$ of affinely spanning points in $\mathbb{K}^{d}$. Let $A$ be the corresponding matrix in the construction of Gale duality for $a_{1},\ldots,a_{n},a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ in $\mathbb{K}^{d}$, and let $B$ be the matrix whose columns are the conjugates of $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$. By the density of $U$ in $S$, there is a sequence $(a_{m})_{m=1}^{\infty}$ of points in $(\mathbb{K}^{d})^{n}$ which are in strong general position and which converge to $a$. For each $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$, let $a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a_{i}^{m}$, and let $A_{m}$ be the matrix whose columns are $(a_{1}^{m},1),\ldots,(a_{n+1}^{m},1)$. As the $a_{m}$ converge to $a$ and the matrices all have full rank, we have that $\ker A_{m}$ converge to $\ker A$ (i.e., if $\pi_{m}$ and $\pi$ are the projections of $\mathbb{K}^{n+1}$ onto $\ker A_{n}$ and $\ker A$, respectively, then $\pi_{m}(x)$ converges to $\pi(x)$ for all $x\in\mathbb{K}^{n+1}$). For each $m$, let $b_{1}^{m},\ldots,b_{n-d}^{m}$ be a basis for $A_{m}$ such that $b_{1}^{m},\ldots,b_{n-d-1}^{m}$ are unit vectors and $b_{n-d}^{m}=(1,\ldots,1,-n)$. Passing to a convergent subsequence if necessary when $11\leq i\leq n-d-$ 1, for each $i\in[n-d-1]$ we have that $b_{i}^{m}$ converges to some $b_{i}^{\prime}$, and that $b_{n-d}^{\prime}=(1,\ldots,1,-n)$. It is easily confirmed that $b_{1}^{\prime},\ldots,b_{n-d-1}^{\prime},b^{\prime}_{n-d}$ is a basis for $\ker A$. While the conjugates of the $b_{i}^{\prime}$ do not necessary correspond to the rows of $\overline{B}$, they do after multiplying $B$ by an appropriate invertible $(n-d)\times(n-d)$ matrix $T$. For each $m$, let $B_{m}$ be the matrix whose rows are $b_{1}^{m},\ldots,b_{n-d}^{m}$. The rows of each $TB_{m}$ are still a basis for $\ker A_{m}$, and the last row of each $TB_{m}$ is still $(1,\ldots,1,-n)$. The matrices $\overline{TB_{m}}$ converge to $\overline{B}$, and so the columns $g_{1}^{m}=(x_{1}^{m},1),\ldots,g_{n-d-1}^{m}=(x_{n-d-1}^{m},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}^{m}$ of the $\overline{TB_{m}}$ converge to $g_{1},\ldots,g_{n+1}$. Thus each $x_{m}=(x_{1}^{m},\ldots,x_{n}^{m})$ lies in $U^{\prime}$ and converges to $x$.
@@
That $U^{\prime}$ is open in $S^{\prime}$ follows similarly. For contradiction, suppose that $U^{\prime}$ were not open in $S^{\prime}$. Thus there is some $x=(x_{1},\ldots,x_{n})$ in $U^{\prime}$ corresponding to some sequence $a=(a_{1},\ldots,a_{n})$ in strong general position and a sequence $(x_{m})_{m=1}^{\infty}$ of points in $S^{\prime}\setminus U^{\prime}$ which converge to $x$. As before, let $A$ the matrix whose columns are $(a,1),\ldots,(a_{n+1},1)$, where $a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$, and let $B$ be the matrix whose columns are $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$. For each $m$, let $B_{m}$ be the matrix whose columns are $g_{1}^{m}=(x_{1}^{m},1),\ldots,g_{n}^{m}=(x_{n}^{m},1),g^{m}_{n+1}=-\sum_{i=1}^{n}g^{m}_{i}$, and choose a basis $b_{1}^{m},\ldots,b_{n-d}^{m}$ for $\ker B_{m}$ so that $b_{n-d}$ is the all 1 vector and the other $b_{i}^{m}$ have norm 1. As $\ker B_{m}$ converges to $\ker B$, after passing to a subsequence if necessary we may assume that each $b_{i}^{m}$ converges to some $b_{i}^{\prime}$, where $b_{n-d}^{\prime}$ is the all one vector. Again, it is easily confirmed that the vectors $b_{1}^{\prime},\ldots,b_{n-d}^{\prime}$ are a basis for $\ker B$. After applying an appropriate invertible transformation to the $b_{i}^{\prime}$ and applying this transform to $b_{1}^{m},\ldots,b_{n-d-1}^{m}$ for all $m$, we may assume that $b_{i}^{\prime}=b_{i}$ for all $1\leq i\leq n-d$. We now let $A_{m}$ be the matrix whose rows are the conjugates of the $b_{i}^{m}$, and we let $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$ be its first $n$ columns excluding the final row of ones. We have $a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ for all $m$ by construction. As the rows of $A$ are the conjugates of the $b_{i}$, $A_{m}$ converges to $A$ and therefore $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$ converges to $a$. On the other hand, each $x_{m}$ corresponds to $a_{m}$, and since $x_{m}$ is in $S^{\prime}\setminus U^{\prime}$, each $a_{m}$ is in $S\setminus U$, thereby contradicting that $U$ is open in $S$. ∎
@@
We now show that the open half-flats of the $r$-fans of Theorems 1.4 and  1.6 must share a large fraction of the points of any typical point set.
@@
Let $d,m\geq 1$ be integers and let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Let $n\geq(r-1)(d+m+1)+1$ if $\mathbb{K}=\mathbb{R}$, let $n\geq(r-1)(2d+m+1)+1$ if $\mathbb{K}=\mathbb{C}$, and suppose that $X$ is a typical set of $n$ points in $\mathbb{K}^{n-d-1}$. Then the $r$-fans $F_{r}=\cup_{j=1}^{r}B_{j}$ guaranteed by Theorem  1.4 and 1.6 satisfy the property that $\sqcup_{j=1}^{r}\operatorname{\mathrm{Int}}(B_{j})$ contains at least $(r-1)(d+1)+1$ points of $X$ when $\mathbb{K}=\mathbb{R}$, and at least $(r-1)(2d+1)+1$ points of $X$ when $\mathbb{K}=\mathbb{C}$.
@@
Suppose that $X=\{x_{1},\ldots,x_{n}\}$ is a typical point set in $\mathbb{K}^{n-d-1}$. Thus $x_{1},\ldots,x_{n}$ corresponds to a sequence of points $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ in strong general position. The distributing $r$-fan $F_{r}=\cup_{j=1}^{r}$ guaranteed by the proofs of Theorem 1.4 and Theorem 1.6 was given by a Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for the linear map $f\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ determined by the $a_{i}$. Moreover, we had $|\operatorname{\mathrm{Int}}(B_{j})\cap X|=|\operatorname{\mathrm{Vert}}(\sigma_{j})|$ for all $j\in[r]$. As the $a_{i}$ are in strong general position, we have $\sum_{j=1}^{r}|\operatorname{\mathrm{Int}}(B_{j})\cap X|=\sum_{j=1}^{r}|\operatorname{\mathrm{Vert}}(\sigma_{j})|\geq(r-1)(d+1)+1$. ∎
@@
We conclude this section by proving the near optimality of the real case of Theorem 1.4 with respect to dimension for all prime powers $r\geq 3$ when $m\geq 2$, and likewise in the complex setting when $r=2$ and $m\geq 1$. For this, it suffices to consider the near optimality of Theorems 1.1 and  1.3, respectively.
@@
    Let $r\geq 3$, let $k\in\{0,\ldots,r-1\}$, and let $n=(r-1)(d+m+1)+k+1$. If $\ell>2+\frac{k}{r-1}$, then there exists a set $X$ of $n+\ell$ affinely spanning points in $\mathbb{R}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equidistributed by any $r$-fan of codimension $r-2$.
@@
    Let $r=2$, let $k\in\{0,1\}$, and let $n=2(d+1)+m+k$. If $\ell>2+\frac{k}{2}$, then there exists a set $X$ of $n+\ell$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equidistributed by any complex regular $2$-fan in $\mathbb{C}^{n-d-1}$.

### v3

@@ 8. Robustness for Typical Point Sets and Optimality of Fan Distributions
We conclude by showing that there is a dense open class of affinely spanning point sets for which the distributing $r$-fans of Theorems 1.4 and 1.6 contain a robust number of points from the given point set. Lastly, we prove various cases of the near optimality of Theorem 1.4 with respect to dimension.
@@ 8.1. Typical Point Sets
Let $a_{1},\ldots,a_{n}$ be a sequence of $n$ ($\mathbb{K}$-) affinely spanning points in $\mathbb{K}^{d}$ and let $a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ be their average. Letting $A$ be the matrix of the Gale transform associated to $a_{1},\ldots,a_{n+1}$, one may choose a basis $b_{1},\ldots,b_{n-d}$ for $\ker A$ with $b_{n-d}=(1,\ldots,1,-n)$, and, for any such choice of such basis, the Gale dual of $a_{1},\ldots,a_{n+1}$ is of the form $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}=-\sum_{i=1}^{n}(x_{i},n)$. We say that the sequence $x_{1},\ldots,x_{n}$ of affinely spanning points in $\mathbb{K}^{n-d-1}$ corresponds to $a_{1},\ldots,a_{n}$. Moreover, by letting $g_{i}=(x_{i},1)$ for $1\leq i\leq n$ and $g_{n+1}=-\sum_{i=1}^{n}g_{i}$, it is not difficult to see that any sequence $x_{1},\ldots,x_{n}$ of affinely spanning points in $\mathbb{K}^{n-d-1}$ corresponds to some sequence $a_{1},\ldots,a_{n}$ of affinely spanning points in $\mathbb{K}^{d}$.
Our condition on affinely spanning point sets $X$ with $n$ elements in $\mathbb{K}^{n-d-1}$ is essentially the Gale dual of a generic condition on affinely spanning point sets $A$ of $n$ elements in $\mathbb{K}^{d}$. We recall that a sequence $a_{1},\ldots,a_{n}$ of $n$ points in $\mathbb{K}^{d}$ is in strong general position [24] if for any $1\leq r\leq n$ and any $r$ pairwise disjoint subsets $I_{1},\ldots,I_{r}$ of $[n]$ one has that the (real) codimension of the intersection of the real affine hulls $\operatorname{\mathrm{Aff}}(A_{j})$ of the $A_{j}=\{a_{i}\mid i\in I_{j}\}$ equals the sum of their respective codimensions, that is,
@@
The central property of strong general position for our purposes is that it immediately implies that any Tverberg $r$-tuple for a linear map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ determined by points $a_{1},\ldots,a_{n}$ in strong general position must satisfy $\sum_{j=1}^{r}|\operatorname{\mathrm{Vert}}(\sigma_{j})|\geq(r-1)(d+1)+1$.
Let $n\geq d+1$. We shall say that an affinely spanning point set $X=\{x_{1},\ldots,x_{n}\}$ in $\mathbb{K}^{n-d-1}$ is typical if $x_{1},\ldots,x_{n}$ corresponds to a sequence $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ in strong general position. While we expect that typical point sets are generic, we shall content ourselves with showing that they form an open and dense subset of the space of all affinely spanning point sets of size $n$ in $\mathbb{K}^{n-d-1}$. Specifically, let $\operatorname{\mathrm{Conf}}_{n}(\mathbb{K}^{n-d-1})=\{(x_{1},\ldots,x_{n})\in(\mathbb{K}^{n-d-1})^{n}\mid x_{i}\neq x_{j}\,\,\text{for all}\,\,i\neq j\}$ be the configuration space of sequences of $n$ pairwise distinct points in $\mathbb{K}^{n-d-1}$. If $S$ denotes the subspace of $(\mathbb{K}^{d})^{n}$ consisting of all affinely-spanning sequences of length $n$ and $U$ is the subset of $S$ consisting of all such sequences which are in strong general position, then the generiticity of strong general position in $\mathbb{K}^{d}$ implies in particular that $U$ is an open dense subset of $S$. An analogous situation holds for typical point sets of size $n$ in $\mathbb{K}^{n-d-1}$.
@@
Let $S^{\prime}$ be the subset $(\mathbb{K}^{n-d-1})^{n}$ consisting of all affinely spanning sequences of length $n$ in $\mathbb{K}^{n-d-1}$ and let $U^{\prime}$ be the subset of $S^{\prime}$ consisting of all such sequences which correspond to a sequence of $n$ points in $\mathbb{K}^{d}$ in strong general position. Then $U^{\prime}\cap\operatorname{\mathrm{Conf}}_{n}(\mathbb{K}^{n-d-1})$ is an open dense subset of $S^{\prime}\cap\operatorname{\mathrm{Conf}}_{n}(\mathbb{K}^{n-d-1})$.
Proposition 8.1 is an immediate consequence of the following “continuity” lemma for the Gale transform.
Lemma 8.2.
Let $n$ and $d$ be positive integers such that $n\geq d+1$. Suppose that $x=(x_{1},\ldots,x_{n})$ is a sequence of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$ which corresponds to a sequence $a=(a_{1},\ldots,a_{n})$ of $n$ affinely spanning points in $\mathbb{K}^{d}$.
  1. (a)
    Suppose that for each $m\geq 1$ there is a sequence $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$ of $n$ affinely spanning points in $\mathbb{K}^{d}$ such that $\lim_{m\to\infty}a_{m}=a$. Then for each $m$ there is a sequence $x_{m}=(x_{1}^{m},\ldots,x_{n}^{m})$ of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$ such that each $x_{m}$ corresponds to $a_{m}$ and $\lim_{m\to\infty}x_{m}=x$.
  2. (b)
    Suppose that for each $m\geq 1$ there is a sequence $x_{m}=(x_{1}^{m},\ldots,x^{m}_{n})$ of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$ such that $\lim_{m\to\infty}x_{m}=x$. Then for each $m$ there is a sequence $a_{m}=(a_{1}^{m},\ldots,a_{m}^{n})$ of $n$ affinely spanning point sets in $\mathbb{K}^{d}$ such that each $x_{m}$ corresponds to $a_{m}$ and $\lim_{m\to\infty}a_{m}=a$.
@@
As $\operatorname{\mathrm{Conf}}_{n}(\mathbb{K}^{n-d-1})$ is open in $(\mathbb{K}^{n-d-1})^{n}$ it suffices to prove that $U^{\prime}$ is an open dense subset of $S^{\prime}$.
We first prove density. Suppose that $x=(x_{1},\ldots,x_{n})$ is a sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$, so that $x$ corresponds to some sequence $a=(a_{1},\ldots,a_{n})$ of affinely spanning points in $\mathbb{K}^{d}$. By the density of $U$ in $S$, there is a sequence $(a_{m})_{m=1}^{\infty}$ of points in $U$ which converges to $a$. By Lemma 8.2(a), there is a sequence $(x_{m})_{m=1}^{\infty}$ of points in $S^{\prime}$ such that each $x_{m}$ corresponds to $a_{m}$ and $x=\lim_{m\to\infty}x_{m}$. As each $a_{m}$ lies in $U$ we have that each $x_{m}$ lies in $U^{\prime}$, which establishes density.
That $U^{\prime}$ is open in $S^{\prime}$ follows similarly. Suppose for a contradiction that $U^{\prime}$ is not open in $S^{\prime}$. Thus there is some $x=(x_{1},\ldots,x_{n})\in U^{\prime}$ corresponding to some $a=(a_{1},\ldots,a_{n})\in U$ and a sequence $(x_{m})_{m=1}^{\infty}$ in $S^{\prime}\setminus U^{\prime}$ which converges to $x$. By Lemma 8.2(b), there is a sequence $(a_{m})_{m=1}^{\infty}$ in $S$ such that $x_{m}$ corresponds to $a_{m}$ and $\lim_{m\to\infty}a_{m}=a$. As each $x_{m}$ lies in $S^{\prime}\setminus U$ we have that each $a_{m}$ lies in $S\setminus U$. This contradicts that $U$ is open in $S$, which completes the proof. ∎
We now prove Lemma 8.2.
Proof of Lemma 8.2.
To prove (a), suppose that $x=(x_{1},\ldots,x_{n})$ is a sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$ which corresponds to some sequence $a=(a_{1},\ldots,a_{n})$ of affinely spanning points in $\mathbb{K}^{d}$. Let $A$ be the $(d+1)\times(n+1)$ matrix in the construction of the Gale dual for $a_{1},\ldots,a_{n},a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ in $\mathbb{K}^{d}$, and let $B$ be the $(n-d)\times(n+1)$ matrix whose columns are the conjugates of $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$. The rows of $B$ form a basis for $\ker A$ by the construction of the Gale transform. For each $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$, let $a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a_{i}^{m}$ and let $A_{m}$ be the matrix in the construction of the Gale dual of $a_{1}^{m},\ldots,a_{n+1}^{m}$.
For each $m$ choose an orthogonal basis $b_{1}^{m},\ldots,b_{n-d}^{m}$ for $\ker A_{m}$ such that $b_{1}^{m},\ldots,b_{n-d-1}^{m}$ are unit vectors and $b_{n-d}^{m}=(1,\ldots,1,-n)$, and let $B_{m}$ be the matrix whose rows are the $b_{i}^{m}$. After passing to convergent subsequence if necessary, we may assume that each $b_{i}^{m}$ converges to some $b_{i}^{\prime}$ with $b_{n-d}^{\prime}=(1,\ldots,1,-n)$. Since the $a_{m}$ converge to $a$ we have that the $A_{m}$ converge to $A$ in the Frobenius norm, from which it is straightforward to show that $b_{1}^{\prime},\ldots,b_{n-d-1}^{\prime},b^{\prime}_{n-d}$ are elements of $\ker A$. By the choice of basis of $\ker A_{m}$ it is likewise readily seen that the $b_{i}^{\prime}$ form a basis for $\ker A$. We let $B^{\prime}$ be the matrix whose rows are the $b_{i}^{\prime}$. Thus $\lim_{m\to\infty}B_{m}=B^{\prime}$.
Now let $T$ be the invertible $(n-d)\times(n-d)$ matrix such that $TB^{\prime}=B$. As $b_{n-d}^{\prime}=b_{n-d}$, the last row of $T$ is the (nd)(n-d)-th standard basis vector $\mathbf{e}_{n-d}\in\mathbb{K}^{n-d}$. For each $m$ the rows of $TB_{m}$ are still a basis for $\ker A_{m}$, and since the last row of $T$ is $\mathbf{e}_{n-d}$ the last row of $TB_{m}$ is still $(1,\ldots,1,-n)$. The Gale dual of $a_{1}^{m},\ldots,a_{n}^{m},a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a^{m}_{i}$ with respect to this basis for $\ker A_{m}$ is therefore of the form $g_{1}^{m}=(x_{1}^{m},1),\ldots,g_{n}^{m}=(x_{n}^{m},1),g^{m}_{n+1}=-\sum_{i=1}^{n}g_{i}^{m}$. Thus each $x_{m}=(x_{1}^{m},\ldots,x_{n}^{m})$ corresponds to $a_{m}$. Since the $B_{m}$ converge to $B^{\prime}$ we have that the $TB_{m}$ converge to $B$ and therefore that $\overline{TB_{m}}$ converges to $\overline{B}$. Thus the $g_{i}^{m}$ converge to $g_{i}$ for each $i\in[n+1]$ and hence the $x_{m}$ converge to $x$.
@@
The proof of part (b) is analogous, where now we reverse the steps in the construction of the Gale dual as in the proof of Proposition 2.1. So suppose that $x=(x_{1},\ldots,x_{n})$ is a sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$ which corresponds to some sequence $a=(a_{1},\ldots,a_{n})$ of affinely spanning points in $\mathbb{K}^{d}$, and moreover that there is a sequence $(x_{m})_{m=1}^{\infty}$ of affinely spanning points in $\mathbb{K}^{d}$ which converges to $x$. Let $A$ be the $(d+1)\times(n+1)$ matrix in the construction for the Gale transform of $a_{1},\ldots,a_{n},a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ and let $B$ be the $(n-d)\times(n+1)$ matrix whose columns are the conjugates of $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$. We have that $\operatorname{\mathrm{Row}}B=\ker A$ by the construction of the Gale transform, so $\ker B=\operatorname{\mathrm{Row}}A$ and thus the rows $w_{1},\ldots,w_{d},w_{d+1}=(1,\ldots,1)$ of $A$ form a basis for $\ker B$. For each $m$, let $B_{m}$ be the matrix whose columns are the conjugates of $g_{1}^{m}=(x_{1}^{m},1),\ldots,g_{n}^{m}=(x_{n}^{m},1),g^{m}_{n+1}=-\sum_{i=1}^{n}g^{m}_{i}$ and choose an orthogonal basis $w_{1}^{m},\ldots,w_{d+1}^{m}$ for $\ker B_{m}$ such that $w^{m}_{d+1}$ is the all 1 vector and the other $w_{i}^{m}$ have norm 1. After passing to a subsequence if necessary we may assume that each $w_{i}^{m}$ converges to some $w_{i}^{\prime}$, where $w_{d+1}^{\prime}$ is the all one vector. As the $x_{m}$ converge to $x$ we have that the $\overline{B_{m}}$ converge to $\overline{B}$ and so that the $B_{m}$ converge to $B$. It follows as before that $w_{1}^{\prime},\ldots,w_{d+1}^{\prime}$ is a basis for $\ker B$.
Now let $A^{\prime}$ be the matrix whose rows are $w_{1}^{\prime},\ldots,w^{\prime}_{d+1}$, and likewise for each $m$ let $A_{m}$ be the matrix whose rows are $w_{1}^{m},\ldots,w_{d+1}^{m}$. We let $T$ be the invertible $(d+1)\times(d+1)$ matrix such that $TA^{\prime}=A$. Since the $w_{i}^{m}$ converge to $w_{i}^{\prime}$, the $A_{m}$ converge to $A^{\prime}$ and so the $TA_{m}$ converge to $A$. For each $m$ the rows of $TA_{m}$ still form a basis for $\ker B_{m}$, and, since $w^{\prime}_{d+1}=w_{d+1}$, the last row of $T$ is $\mathbf{e}_{d+1}$ and so the last row of $TA_{m}$ is still the all one vector. Finally, for each $m$ let $a_{1}^{m},\ldots,a_{n+1}^{m}$ be the columns of the matrix formed by removing the final row from $TA_{m}$ and let $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$. The proof of Proposition 2.1 shows that $g_{1}^{m},\ldots,g_{n+1}^{m}$ is the Gale dual of $a_{1}^{m},\ldots,a_{n+1}^{m}$ obtained by letting the rows of $B_{m}$ serve as the chosen basis for $\ker TA_{m}$. Because the last row of $B_{m}$ is $(1,\ldots,1,-n)$ we have that $u_{n+1}=\frac{1}{n}\sum_{i=1}^{n}u_{i}$ for each $u=(u_{1},\ldots,u_{n+1})\in\ker B_{m}$ and therefore that $a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a^{m}_{i}$. Thus each $x_{m}$ corresponds to $a_{m}$. Since the $TA_{m}$ converge to $A$ we have that the $a_{m}$ converge to $a$. ∎
@@
Next, we show that for a typical point set the open half-flats of the $r$-fans of Theorems 1.4 and  1.6 contain a sizable number of points.
@@
Let $d,m\geq 1$ be integers and let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Let $n\geq(r-1)(d+m+1)+1$ if $\mathbb{K}=\mathbb{R}$, let $n\geq(r-1)(2d+m+1)+1$ if $\mathbb{K}=\mathbb{C}$, and suppose that $X$ is a typical set of $n$ points in $\mathbb{K}^{n-d-1}$. Then the $r$-fans $F_{r}=\cup_{j=1}^{r}B_{j}$ guaranteed by Theorem  1.4 and 1.6 satisfy the property that $\sqcup_{j=1}^{r}\operatorname{\mathrm{Int}}(B_{j})$ contains at least (r1)(d+1)+1(r-1)(d+1)+1 points of $X$ when $\mathbb{K}=\mathbb{R}$, and at least (r1)(2d+1)+1(r-1)(2d+1)+1 points of $X$ when $\mathbb{K}=\mathbb{C}$.
@@
Suppose that $X=\{x_{1},\ldots,x_{n}\}$ is a typical point set in $\mathbb{K}^{n-d-1}$. Thus $x_{1},\ldots,x_{n}$ corresponds to a sequence of points $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ in strong general position. The distributing $r$-fan $F_{r}=\cup_{j=1}^{r}$ guaranteed by the proofs of Theorem 1.4 and Theorem 1.6 was given by a Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for the linear map $f\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ determined by the $a_{i}$. Moreover, we had $|\operatorname{\mathrm{Int}}(B_{j})\cap X|=|\operatorname{\mathrm{Vert}}(\sigma_{j})|$ for all $j\in[r]$. As the $a_{i}$ are in strong general position, we have $\sum_{j=1}^{r}|\operatorname{\mathrm{Int}}(B_{j})\cap X|=\sum_{j=1}^{r}|\operatorname{\mathrm{Vert}}(\sigma_{j})|\geq(r-1)(d+1)+1$ when $\mathbb{K}=\mathbb{R}$ and $\sum_{j=1}^{r}|\operatorname{\mathrm{Int}}(B_{j})\cap X|\geq(r-1)(2d+1)+1$ when $\mathbb{K}=\mathbb{C}$. ∎
@@
Finally, we prove the near optimality of the real case of Theorem 1.4 with respect to dimension for all prime powers $r\geq 3$ when $m\geq 2$, and likewise in the complex setting when $r=2$ and $m\geq 1$. For this, it suffices to consider the near optimality of Theorems 1.1 and  1.3, respectively.
@@
    Let $r\geq 3$, let $k\geq 0$, and let n=(r1)(d+m+1)+k+1n=(r-1)(d+m+1)+k+1. If $\ell>2+\frac{k+2}{r-2}$, then there exists a set $X$ of $n+\ell$ affinely spanning points in $\mathbb{R}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equidistributed by any conical $r$-fan.
@@
    Let $k\geq 0$ and let $n=2(d+1)+m+k$. If $\ell>k+3$ then there exists a set $X$ of $n+\ell$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equidistributed by any complex regular $2$-fan in $\mathbb{C}^{n-d-1}$.